OpenStudy (anonymous):

A bag contains 2 steel balls and 5 brass balls. The total weight is 13 pounds. If 2 steel balls are added and 2 brass balls are removed, the weight decreases to 12 pounds. How much does each kind of ball weigh?

OpenStudy (anonymous):

steel, 4 pounds; brass, 1 pound steel, 1.5 pounds; brass, 2 pounds steel, 0.75 pounds; brass, 3 pounds steel, 3 pounds; brass, 4 pounds

OpenStudy (mathstudent55):

You can set up a system of equations to solve it.

OpenStudy (anonymous):

Ok..

OpenStudy (mathstudent55):

You need to let the variables represent the weights of the balls since that is what we are trying to find. Let s = weight of a steel ball Let b = weight of a brass ball

OpenStudy (anonymous):

Ok so would it be something like s+b=13

OpenStudy (mathstudent55):

At first, we are told about 2 steel balls and 5 brass balls. 2 steel balls weigh 2s 5 brass balls weigh 5b Their total weight is 13 That means we can write this equation: 2s + 5b = 13

OpenStudy (mathstudent55):

You're on the right track, but remember it's 2 steel balls and 5 brass balls. That is why it's 2s + 5b on the left side.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I believe your answer would be Steel : 1.5 pounds and Brass: 2 pounds Like mathstudent55 said you can use a system of equations. Let S= Steel balls and B= Brass balls 2S+5B=13 4S+3B=12 (This is the second one 2 Steel balls are added and 2 brass balls are removed) You want to isolate a variable and in this case you can do so if you multiply the first equation by 2 so that you can do 4S - 4S leaving you with one variable. 2(2S+5B)= 13(2) 4S+10B=26 Now you can do 4S+10B=26 - 4S+3B=12 7B=14 B=2 This means that the brass balls weigh 2 pounds. Now if the brass balls weigh 2 pounds then plug it back into the original equation (either one) 2S+5(2)=13 2S+10=13 2S=3 S=1.5 That is your weight for a Steel ball so your answer is the second choice.

OpenStudy (mathstudent55):

The second time, there are 2 steel balls extra and 2 brass balls fewer, so we get 4s + 3b We are told the weight is now 12 lb, so we get the equation: 4s + 3b = 12

OpenStudy (anonymous):

@WordGEEK Thats exactly what i just got.

OpenStudy (mathstudent55):

That means we have this system of equations: 2s + 5b = 13 4s + 3b = 12

OpenStudy (mathstudent55):

Now we need to solve the system of equations. There are several methods for doing this, but in this case, addition seems to be easy to do.

OpenStudy (mathstudent55):

Multiply the entire first equation by -2 on both sides. Then write down the original second equation below it. -4s - 10b = -26 4s + 3b = 12 Now add the equations. The s terms will add to zero leaving only b. 0 - 7b = -14 -7b = -14 Divide both sides by -2: b = 2 A brass ball weighs 2 lb

OpenStudy (mathstudent55):

Now you can use either of the two original equations to find s, using the value of b we already know, 2. 2s + 5b = 13 2s + 5(2) = 13 2s + 10 = 13 2s = 3 s = 1.5 A steel ball weighs 1.5 lb The answer is: a steel ball weighs 1.5 lb, and a brass ball weighs 2 lb

OpenStudy (anonymous):

Thanks @mathstudent55 and @WordGEEK

OpenStudy (mathstudent55):

You're welcome.